### v1

Abstract
Fejér monotonicity is a well-established property commonly observed in sequences generated by optimization algorithms. In this paper, we introduce an extension of this property, called Fejér* monotonicity, which was initially proposed in [SIAM J. Optim., 34(3), 2535–2556 (2024)]. We discuss and build upon the concept by exploring its behavior within Hilbert spaces, presenting an illustrative example and insightful results regarding weak and strong convergence. We also compare Fejér* monotonicity with other weak notions of Fejér-like monotonicity, to better establish the role of Fejér* monotonicity in optimization algorithms.
Keywords:
Fejér monotonicity convex analysis optimization algorithms
MSC:
49M27 65K05 65B99 90C25
journal: JOTA
1 Introduction
A useful concept for analyzing the convergence of optimization algorithms is the notion of Fejér monotonicity. A sequence $(x^{k})_{k\in\mathds{N}}$ defined in a Hilbert space $\mathcal{H}$ is said to be Fejér monotone with respect to a nonempty set $M\subset\mathcal{H}$ if for all $x\in M$ and all $k\in\mathds{N}$ we have $\lVert x^{k+1}-x\rVert\leq\lVert x^{k}-x\rVert$. It is well-known that if $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone with respect to $M$, then $(x^{k})_{k\in\mathds{N}}$ is bounded, and if all its weak cluster points belong to $M$, then the entire sequence $(x^{k})_{k\in\mathds{N}}$ converges weakly to some point lying in $M$; see, for instance, (Bauschke:1996, , Theorem 2.16), and Combettes:2009 for further details and historical notes.
The notion of Fejér monotonicity is named after Hungarian mathematician Lipót Fejér (1880-1959) who introduced this notion in Fejer:1922 and further developed in Fejer:1937 and later in Agmon:1954 ; Motzkin:1954 . Fejér was the force behind the highly successful Hungarian school of analysis. His influence on the development of mathematics in the last century should be acknowledged. Indeed, Fejér was the advisor of mathematicians such as John von Neumann, Paul Erdös, George Pólya, Pál Turán and Gábor Szegö.
Weaker notions of Fejér monotonicity have been considered in the literature. For instance, the concept of quasi-Fejér monotonicity, first presented in Ermolev:1969 and further elaborated in several works such as Iusem:1994 ; Combettes:2001a .
In this work, we explore a slightly different notion, called Fejér* monotonicity. It was first introduced inBehling:2024c as a key ingredient to show that the Circumcentered-Reflection Method (CRM) Behling:2024b , proposed by these authors, solves the Convex Feasibility Problem (CFP). The main difference with the usual Fejér monotonicity notion relies on the fact that now the nonincreasing distance property holds for the tail of the sequence, starting at some index which depends on the considered point $x\in M$.
This paper is organized as follows. We start in Section 2 with a brief discussion on the role of Fejér monotonicity in optimization algorithms. In Section 3 we present some basic properties of Fejér* monotonicity, and we make a discussion, upon an illustrative example on properties of Fejér* monotonicity. In Section 4 we compare Fejér* sequences and quasi-Fejér sequences. Finally, in Section 5 we present some concluding remarks.
2 Fejér* monotonicity in optimization algorithms
First, let us formally introduce the notion of Fejér* monotonicity.
Definition 1 (Fejér* monotonicity)
Let $M\subset\mathcal{H}$ be a nonempty set and consider a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$. We say that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone with respect to $M$ if for any point $x\in M$, there exists $N(x)\in\mathds{N}$ such that, for all $k\geq N(x)$,
$\left\lVert x^{k+1}-x\right\rVert\leq\left\lVert x^{k}-x\right\rVert.$
We address now the role of Fejér and Fejér monotonicity in connection with fixed-point and feasibility problems. Consider an operator $T:\mathds{R}^{n}\to\mathds{R}^{n}$. We remind that $T$ is a contraction if there exists $\tau\in(0,1)$ such that
$\left\lVert T(x)-T(y)\right\rVert\leq\tau\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
It is well known, and elementary, that a contraction has a unique fixed point (i.e., a point $\bar{x}\in\mathds{R}^{n}$ such that $T(\bar{x})=\bar{x}$) and that the sequence obtained by iterating the operator (i.e., the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{n}$ defined as $x^{k+1}=T(x^{k})$ starting at any $x^{0}\in\mathds{R}^{n}$), converges to such unique fixed point.
The situation changes when we consider notions which are weaker than the contraction property. The most basic one is that of nonexpansiveness. An operator $T$ is nonexpansive if
$\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
for all $x,y\in\mathds{R}^{n}$.
The set of fixed points of a nonexpansive operator $T$ may be empty, or infinite, and the behavior of the sequence $(x^{k})_{k\in\mathds{N}}$ obtained by iterating $T$ may fail to converge, but given any fixed point $\bar{x}$ of $T$, it follows immediately from Section 2 that $\left\lVert x^{k+1}-\bar{x}\right\rVert\leq\left\lVert x^{k}-\bar{x}\right\rVert$, i.e., the sequence is Fejér monotone to the set $M$ of fixed points of $T$, and thus it enjoys the basic properties of Fejér monotone sequences, namely that it is bounded, and that if some of its cluster points belongs to $M$, then the sequence $(x^{k})_{k\in\mathds{N}}$ converges to it; see, for instance, Bauschke:2017a for more on nonexpansiveness and Fejér monotonicity.
Among the nonexpansive operators, projections onto convex sets are quite significant. We remind that, given a closed and convex set $C\subset\mathds{R}^{n}$, the projection onto $C$, $P_{C}:\mathds{R}^{n}\to C$, is defined as $P_{C}(x)\coloneqq\operatorname*{arg\,min}_{y\in C}\left\lVert x-y\right\rVert$. Convex combinations of projections (i.e., operators of the form $\bar{P}=\sum_{i+1}^{m}\alpha_{i}P_{C_{i}}$ with $\alpha_{i}\in[0,1],\sum_{i=1}^{m}\alpha_{i}=1$, and $C_{i}\subset\mathds{R}^{n}$ convex) are nonexpansive, and the same holds for compositions of projections, i.e., operator of the form $\widehat{P}=P_{C_{m}}\circ\dots\circ P_{C_{1}}$. The set of fixed points of a projection $P_{C}$ is $C$, and the sets of fixed points of $\bar{P}$ and $\widehat{P}$, as defined above are equal to the set $C:=\cap_{i=1}^{m}C_{i}$. Iterating the operators $\bar{P}$ and $\widehat{P}$, one obtains the so-called Simultaneous Projections and Sequential Projections algorithms, which aim to solve the (CFP), consisting of finding a point in the intersection $C$ of closed and convex sets $C_{1},\dots,C_{m}$. See Bauschke:1996 for a detailed discussion on these methods, which can be traced back to the seminal works of Cimmino Cimmino:1938 and Kaczmark Kaczmarz:1937 . In view of the discussion above, the sequences generated by these two methods are Fejér monotone to $C$.
In order to show that the sequences generated by these algorithms do converge to a point in $C$, one needs a bit more than nonexpansiveness. Two suitable properties, slightly stronger than nonexpansiveness but weaker than the contraction property, are
  1. (i)
    firm nonexpansiveness, meaning that
    $\left\lVert T(x)-T(y)\right\rVert^{2}\leq\left\lVert x-y\right\rVert^{2}-\left\lVert(T(x)-T(y))-(x-y)\right\rVert^{2}$
    for all $x,y\in\mathds{R}^{n}$; and
  2. (ii)
    nonexpansiveness plus, meaning that
    $\left\lVert T(x)-T(y)\right\rVert\leq\left\lVert x-y\right\rVert$
    for all $x,y\in\mathds{R}^{n}$, and $T(x)-T(y)=x-y$ whenever equality holds in (ii).
Projections onto convex sets, as well as the above defined operators $\bar{P}$ and $\widehat{P}$, are firmly nonexpansive. Convex combinations of firmly nonexpansive operators may fail to be firmly nonexpansive, but they are nonexpansive plus (see Arefidamghani:2023 ). With the help of the Fejér convergence properties and the notion of firm nonexpansiveness, it is easy to prove that the Simultaneous and the Successive Projection algorithms do converge to a point in $C$, whenever $C$ is nonempty (see Bauschke:1996 ).
One way to accelerate these methods for CFP is to approximate the set $C$ from the inside, i.e., to replace at iteration $k$ the set $C$ by a set $C^{k}$ such that
$C^{k}\subset C^{k+1}\subset\dots\subset C,$
for all $k$. For instance, if $C\coloneqq\{x\in\mathds{R}^{n}\mid g(x)\leq 0\}$ for some convex $g:\mathds{R}^{n}\to\mathds{R}$, then $C^{k}$ could be chosen as
$C^{k}\coloneqq\{x\in\mathds{R}^{n}\mid g(x)+\epsilon_{k}\leq 0\},$
where $(\epsilon_{k})_{k\in\mathds{N}}$ is a sequence of positive real numbers decreasing to 00. Then one could consider the sequence $(x^{k})_{k\in\mathds{N}}$ given by $x^{k+1}=T_{k}(x^{k})$, where $T_{k}$ is the projection onto a convex set which contains $C^{k}$. In such a case, the nonexpansiveness of $T^{k}$ gives $\left\lVert x^{k+1}-y\right\rVert\leq\left\lVert x^{k}-y\right\rVert$ but only for points $y\in C^{k}$, not in the whole set $C$.
However, if we take $C^{k}$ as in Section 2, and $C$ has nonempty interior, any point $y\in$ int$(C)$ (i.e., such that $g(y)<0$), will belong to $C^{k}$ for all $k$ bigger than $k(y)$, defined as the first integer such that $\epsilon_{k(y)}<-g(y)$, due to Section 2, but this is exactly the same as saying that $(x^{k})_{k\in\mathds{N}}$ is Fejér monotone to int$(C)$.
Hence, Fejér* monotonicity is precisely what is needed for the analysis of algorithms which use approximation schemes as in Section 2. Among them, we mention the algorithms in Iusem:1986a ; Iusem:1986 for CFP. These methods employ the scheme in Section 2 combined with operators related to $\bar{P}$ and $\widehat{P}$, where the projections onto the sets $C_{i}$, approximated from the inside by sets $C^{k}_{i}$, as in Section 2, are replaced by projections onto hyperplanes separating $x^{k}$ from $C^{k}_{i}$. In these cases, the Fejér monotonicity ensures convergence to a point lying possibly in the boundary of $C$ (despite the fact that the sequence is Fejér monotone to int$(C)$). A similar situation occurs with the algorithm in Behling:2024c , where the algorithm in Iusem:1986a is enhanced through the addition of a circumcentering step, using the machinery developed during the last few years in Araujo:2022 ; Arefidamghani:2021 ; Arefidamghani:2023 ; Behling:2018 ; Behling:2018a ; Behling:2020 ; Behling:2021b ; Behling:2023 ; Behling:2024 ; Behling:2024b ; Behling:2024c .
It is worthwhile to note that in the case of Iusem:1986 ; Iusem:1986a ; Behling:2024c , under a Slater hypothesis (i.e., when $\operatorname{int}(C)$ is nonempty), if the $\epsilon_{k}$’s decrease to 00 slowly enough, e.g., so that $\sum_{k=0}^{\infty}\epsilon_{k}=+\infty$, for instance $\epsilon_{k}=1/k$, convergence is indeed finite, i.e., there exists $\bar{k}$ such that $x^{\bar{k}}$ belongs to $C=\bigcap_{i=1}^{m}C_{i}$.
3 Properties of Fejér* monotonicity
The first result of this section was already proved in (Behling:2024c, , Prop. 2.5). However, for the sake of completeness, we present it here.
Proposition 1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
  1. (i)
    $(x^{k})_{k\in\mathds{N}}$ is bounded;
  2. (ii)
    for every $x\in M$, the scalar sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ converges;
  3. (iii)
    $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$.
Proof
For proving (i), take any point $x\in M$. From the definition of Fejér* monotonicity, we conclude that $x^{k}$ belongs to the ball with center at $x$ and radius $\left\lVert x^{N(x)}-x\right\rVert$ for all $k\geq N(x)$. Consequently, $(x^{k})_{k\in\mathds{N}}$ is bounded.
Item (ii) is a direct consequence of Definition 1, since the $N(x)$-tail of sequence $(\left\lVert x^{k}-x\right\rVert)_{k\in\mathds{N}}$ is monotone and bounded, so the sequence converges.
For item (iii), take any $w\in\operatorname{conv}(M)$. Thus, $w$ can be written as $w=\sum_{i=1}^{p}\lambda_{i}w^{i}$, where $w^{i}\in M$, $\lambda_{i}\in[0,1]$ for all $i=1,2,\ldots,p$, $\sum_{i=1}^{p}\lambda_{i}=1$ and $p\in\mathds{N}$. Taking into account that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$, for each $i=1,2,\ldots,p$, there exists $N(w^{i})$ such that $\left\lVert x^{k+1}-w^{i}\right\rVert\leq\left\lVert x^{k}-w^{i}\right\rVert$ for all $k\geq N(w^{i})$. Squaring both sides of the last inequality and rearranging the terms, we get
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w^{i}\right\rangle.$ (1)
The above inequality also holds for all $k\geq K\coloneqq\max\{N(w^{1}),N(w^{2}),\ldots,N(w^{p})\}$. Multiplying Eq. 1, for each $i=1,2,\ldots,p$, by the respective $\lambda_{i}$ and adding up, we obtain
00 0 $\leq\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},\sum_{i=1}^{p}\lambda_{i}w^{i}\right\rangle$ (2)
$=\left\lVert{x^{k}}\right\rVert^{2}-\left\lVert{x^{k+1}}\right\rVert^{2}-2\left\langle x^{k}-x^{k+1},w\right\rangle,$ (3)
and therefore, $\left\lVert x^{k+1}-w\right\rVert\leq\left\lVert x^{k}-w\right\rVert$, for all $k\geq K$. So $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$, as the proof is complete. ∎
We remark that a Fejér monotone sequence with respect to a nonempty set $M\subset\mathcal{H}$ is also Fejér monotone with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, the closure of $\operatorname{conv}(M)$; see (Bauschke:2023, , Lem. 2.1(ii)). Furthermore, Proposition 1(iii) guarantees that a Fejér* monotone sequence w.r.t. $M$ is as well Fejér* monotone w.r.t. $\operatorname{conv}(M)$. However, the Fejér* monotonicity cannot always be extended to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, as shown in the next example.
Example 1
Let $M\coloneqq(w^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ with $w^{k}\coloneqq(1/2^{k},0)$, for all $k\in\mathds{N}$. Define the sequence $(x^{k})_{k\in\mathds{N}}\subset\mathds{R}^{2}$ by $x^{0}\coloneqq(0,2)$ and, for $\ell\in\mathds{N}$,
$x^{2\ell+1}\coloneqq x^{2\ell}+(1/2^{\ell},0)\;,\quad x^{2\ell+2}\coloneqq\left(0,\sqrt{\|x^{2\ell+1}-(1,0)\|^{2}-1}\right).$
If $k$ is even, $x^{k+1}$ is obtained by moving $1/2^{(k/2)}$ horizontally to the right. If $k$ is odd, $x^{k+1}$ is the intersection between the vertical axis and the arc centered in $w^{0}=(1,0)$ passing through $x^{k}$; see Figure 1.
Figure 1: From Example 1: set $M$, first nine points of sequence $(x^{k})_{k\in\mathds{N}}$, $\operatorname{conv}(M)=(0,1]\times\{0\}$, $\operatorname{cl}\left({\operatorname{conv}(M)}\right)\setminus\operatorname{conv}(M)=\{(0,0)\}$ and limit point $x^{*}$ of sequence $(x^{k})_{k\in\mathds{N}}$.
We now show that $(x^{k})_{k\in\mathds{N}}$ is Fejér* with respect to $M$. Denote $x^{k}=(\alpha_{k},\beta_{k})$, $k\in\mathds{N}$. One can see that $1\leq\beta_{k+1}\leq\beta_{k}$ as well as $\|x^{k}-w^{0}\|\geq 1$, for all $k\in\mathds{N}$, so that the sequence $(x^{k})_{k\in\mathds{N}}$ is well-defined.
Now, consider $\lambda\in(0,1]$ and set $w\coloneqq(\lambda,0)\in\operatorname{conv}(M)$. It so happens that $\|x^{k+1}-w\|\leq\|x^{k}-w\|$, for all odd $k$. Indeed,
$\beta_{k+1}=\sqrt{\alpha_{k}^{2}-2\alpha_{k}+\beta_{k}^{2}}\leq\sqrt{\alpha_{k}^{2}-2\alpha_{k}\lambda+\beta_{k}^{2}}\eqqcolon\gamma_{k},$
and $\gamma_{k}$ is such that $z^{k}\coloneqq(0,\gamma_{k})$ belongs to the circumference centered in $w$ and passing through $x^{k}$. In other words, $z^{k}$ complies with $\|z^{k}-w\|=\|x^{k}-w\|$. Thus,
$\|x^{k+1}-w\|=\sqrt{\beta_{k+1}^{2}+\lambda^{2}}\leq\sqrt{\gamma_{k}^{2}+\lambda^{2}}=\|x^{k}-w\|.$
On the other hand, for $w^{\ell}=(1/2^{\ell},0)\in{M}$ with $\ell\geq 0$, consider $N(w^{\ell})=2\ell$. For all even $k\geq N(w^{\ell})$, we have $\alpha_{k}=0$, $\alpha_{k+1}\leq 1/2^{\ell}$ and $\beta_{k+1}=\beta_{k}$. So, $\|x^{k+1}-w^{\ell}\|\leq\|x^{k}-w^{\ell}\|$ and the sequence is Fejér* monotone w.r.t. $M$.
Note that $\operatorname{conv}(M)=(0,1]\times\{0\}$ and also it holds that $\bar{w}=(0,0)\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. We show now that $(x^{k})_{k\in\mathds{N}}$ is not Fejér* monotone w.r.t. $\{(0,0)\}$. Indeed, for $k=2\ell$, $\ell\in\mathds{N}$, we have $x^{2\ell}=(0,\beta_{2\ell})$ and $x^{{2\ell}+1}=(1/2^{\ell},\beta_{{2\ell}+1})=(1/2^{\ell},\beta_{2\ell})$. Thus,
$\|x^{{2\ell}+1}-\bar{w}\|^{2}=\dfrac{1}{2^{2\ell}}+\beta_{2\ell}^{2}>\beta_{2\ell}^{2}=\|x^{2\ell}-\bar{w}\|^{2}.$
Finally, just for the sake of completion, we show that sequence $(x^{k})_{k\in\mathds{N}}$ converges to $x^{*}\coloneqq(0,\sqrt{4/3})$. In fact, for $k=2\ell$, $\ell\in\mathds{N}$, we have from Example 1 that
$\beta_{2\ell+2}^{2}$ $=\alpha_{2\ell+1}^{2}-2\alpha_{2\ell+1}+\beta_{2\ell+1}^{2}=\dfrac{1}{2^{2\ell}}-2\dfrac{1}{2^{\ell}}+\beta_{2\ell}^{2}$ (4)
$=\displaystyle\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{2j}}-2\sum\limits_{j=0}^{\ell}\dfrac{1}{2^{j}}+\beta_{0}^{2}=\displaystyle\sum_{j=0}^{\ell}\dfrac{1}{2^{2j}}\stackrel{{\scriptstyle\ell\to+\infty}}{{\to}}\dfrac{4}{3}.$ (5)
Clearly, we have $\left\lVert x^{2\ell+1}-x^{2\ell}\right\rVert\to 0$ when $\ell\to+\infty$, so $x^{k}\to x^{*}$, as desired. ∎
Remark 1
One might argue that the set $M$ in Example 1 is somehow special, for $M$ has empty interior. However, it is possible to prove the same result for an extended example where the underlying set has nonempty interior, say, for instance, $M\coloneqq\{(w^{\prime},w^{\prime\prime})\in\mathds{R}^{2}\mid 0<w^{\prime}\leq 1,-1<w^{\prime\prime}\leq 0\}$.
In the next proposition the closure of the convex hull of the underlying set plays a role. Even though a Fejér* monotone sequence w.r.t. $M\subset\mathcal{H}$ is not Fejér* monotone w.r.t. $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, one can characterize Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull. This result can be useful for the analysis regarding weak convergence of optimization algorithms that generate Fejér* monotone sequences.
Proposition 2
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then,
  1. (i)
    for every $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ the scalar sequence $(\left\lVert x^{k}-\bar{x}\right\rVert)_{k\in\mathds{N}}$ converges;
  2. (ii)
    for every $\bar{x}_{1},\bar{x}_{2}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, it holds that the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges.
Proof
Regarding item (i), recall that Proposition 1(iii) implies that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $\operatorname{conv}(M)$. Thus, for any $w\in\operatorname{conv}(M)$, Proposition 1(ii) yields that the scalar sequence $(\left\lVert x^{k}-w\right\rVert)_{k\in\mathds{N}}$ converges.
Now, take $\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, and consider a sequence $(w^{\ell})_{k\in\mathds{N}}\subset\operatorname{conv}(M)$ such that $w^{\ell}\rightarrow\bar{x}$. Using the triangle inequality, we have, for all $\ell\in\mathds{N}$,
$-\left\|w^{\ell}-\bar{x}\right\|\leq\left\|x^{k}-\bar{x}\right\|-\left\|x^{k}-w^{\ell}\right\|\leq\left\|w^{\ell}-\bar{x}\right\|.$ (6)
[…]
$-\left\|w^{\ell}-\bar{x}\right\|$ $\leq\liminf_{k}\left\|x^{k}-\bar{x}\right\|-\lim_{k}\left\|x^{k}-w^{\ell}\right\|$ (7)
$\leq\limsup_{k}\left\lVert x^{k}-\bar{x}\right\rVert-\lim_{k}\|x^{k}-w^{\ell}\|\leq\left\|w^{\ell}-\bar{x}\right\|,$ (8)
because, as aforementioned, item Proposition 1(ii) says that, for every $\ell\in\mathds{N}$, $\lim_{k}\left\|w^{\ell}-x^{k}\right\|$ exists. Now, as $\ell$ goes to infinity, we have $\left\|w^{\ell}-\bar{x}\right\|\to 0$, and we conclude that
$\liminf_{k}\left\|x^{k}-\bar{x}\right\|=\limsup_{k}\left\|x^{k}-\bar{x}\right\|=\lim_{k}\left\|x^{k}-\bar{x}\right\|,$
so that we have the desired result.
As for item (ii), let $\bar{x}_{1},\bar{x}_{2}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$. Then,
$\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle$ $=\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}-\bar{x}_{1}\right\rangle+\left\langle\bar{x}_{1}-\bar{x}_{2},\bar{x}_{1}\right\rangle$ (9)
$=\frac{1}{2}\left(\left\|x^{k}-\bar{x}_{2}\right\|^{2}-\left\|\bar{x}_{1}-\bar{x}_{2}\right\|^{2}-\left\|x^{k}-\bar{x}_{1}\right\|^{2}\right)+\left\langle\bar{x}_{1}-\bar{x}_{2},\bar{x}_{1}\right\rangle.$ (10)
Now, item (i) implies that the sequences $(\left\|x^{k}-\bar{x}_{1}\right\|)_{k\in\mathds{N}}$ and $(\left\|x^{k}-\bar{x}_{2}\right\|)_{k\in\mathds{N}}$ converge. Therefore, the sequence $\left(\left\langle\bar{x}_{1}-\bar{x}_{2},x^{k}\right\rangle\right)_{k\in\mathds{N}}$ converges. ∎
4 Fejér* and quasi-Fejér sequences
In this section we discuss Fejér* monotonicity relation with quasi-Fejér monotone sequences (as addressed in Combettes:2001a ). First, we establish the following notation: $\ell_{+}$ is the set of real sequences with nonnegative entries, and $\ell^{1}$ is the set of real sequences with finite $\ell^{1}$-norm.
Definition 2 (quasi-Fejér monotonicity (Combettes:2001a, , Def. 1.1))
Let $M\subset\mathcal{H}$ be a nonempty set. We say that a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, with respect to $M$, is
  1. (i)
    quasi-Fejér of Type I if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|\leq\left\|x^{k}-x\right\|+\varepsilon_{k};$
  2. (ii)
    quasi-Fejér of Type II if there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $x\in M$ and $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k};$
  3. (iii)
    quasi-Fejér of Type III if for all $x\in M$ there exists a sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ such that, for all $k\in\mathds{N}$, it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}+\varepsilon_{k}.$
Sure enough, Fejér sequences are quasi-Fejér of Type I sequences. Also, one can clearly see that Type II implies Type III. Moreover, it was established in (Combettes:2001a, , Prop. 3.2) that if $M$ is bounded, then Type I implies Type II.
Quasi-Fejér of Type III sequences are the most general ones. However, due to this generality, they are hardly used for the analysis of optimization algorithms. In fact, the most common type of quasi-Fejér sequences used in the literature are quasi-Fejér of Type II sequences; see, for instance, Alber:1998 ; Iusem:1994 .
The next theorem shows that Fejér* monotonicity sequences are also quasi-Fejér of Type III.
Theorem 4.1
Let $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$ be a Fejér* monotone sequence w.r.t. a nonempty set $M$ in $\mathcal{H}$. Then, $(x^{k})_{k\in\mathds{N}}$ is a quasi-Fejér of Type III sequence w.r.t. $M$.
Proof
Suppose that $(x^{k})_{k\in\mathds{N}}$ is Fejér* monotone w.r.t. $M$. Then, in view of Definition 1, for any $x\in M$, there exists $N(x)$ such that, for all $k\geq N(x)$, it holds that
$\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}.$
Let us now define the sequence $(\varepsilon_{k})_{k\in\mathds{N}}$. For $k\geq N(x)$, we set $\varepsilon_{k}\coloneqq 0$. On the other hand, for $k<N(x)$, we might have
$\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|,$
so we define
$\varepsilon_{k}\coloneqq\begin{cases*}\left\|x^{k+1}-x\right\|^{2}-\left\|x^{k}-x\right\|^{2},&\text{if $\left\|x^{k+1}-x\right\|>\left\|x^{k}-x\right\|$},\\ 0,&\text{otherwise}.\end{cases*}$
Thus, it holds true that sequence $(\varepsilon_{k})_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$ and therefore sequence $(x^{k})_{k\in\mathds{N}}$ is quasi-Fejér of Type III w.r.t. $M$. ∎
We note that the concepts of Fejér* and quasi-Fejér of Type III sequences clearly are not equivalent, with the latter being more general than the former.
Actually, if $M$ is a singleton, then any Fejér* monotone sequence is quasi-Fejér of Type II, using the same arguments used in the proof of Theorem 4.1. Of course, in this case one can easily build quasi-Fejér sequences of Type II that are not Fejér*. Nevertheless, if a sequence is quasi-Fejér of Type II w.r.t. a general $M$, then it is Type II with respect to $\operatorname{cl}\left({\operatorname{conv}(M)}\right)$ (see (Combettes:2001a, , Prop. 3.4)). Hence, in view of Example 1, we have a Fejér* sequence that it is not quasi-Fejér of Type II monotone. Note that we can construct a non-singleton set $M$, for instance, $M\coloneqq[-1,0]\times\{0\}\subset\mathds{R}^{2}$, such that the same sequence $(x^{k})_{k\in\mathds{N}}$ of Example 1 is quasi-Fejér of Type II, but not Fejér* monotone, w.r.t. $M$. More than not being equivalent, we have just seen that the concepts of Fejér* and quasi-Fejér of Type II are not contained in each other, that is, those ideas are independent.
The previous theorem reveals a particular quasi-Fejér of type III sequence that is related to optimization algorithms, through Fejér* monotonicity Behling:2024c . Thus, we end our discussion by presenting some characterizations of Fejér* monotonicity in terms of the weak convergence of the underlying sequence. These statements are inspired by Combettes:2001a and could be derived from the results therein, since Fejér* monotonicity implies quasi-Fejér of type III monotonicity. However, we present them here, for the sake of completeness. Moreover, the provided proofs are simpler and more clear for the reader, in the context of Fejér* monotonicity.
From now on, for a sequence $(x^{k})_{k\in\mathds{N}}\subset\mathcal{H}$, we denote by $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its weak cluster points and by $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ the set of its strong cluster points. First, we give portrayals of Fejér* monotonicity in terms of the weak convergence of the sequence.
Proposition 3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to $M\subset\mathcal{H}$. Then,
  1. (i)
    $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\neq\emptyset$.
  2. (ii)
    for any $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\};$
  3. (iii)
    if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$, then $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly;
  4. (iv)
    if $x^{k}\rightharpoonup\bar{x}\in\operatorname{cl}\left({\operatorname{conv}(M)}\right)$, then the scalar sequence $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$ converges for every $y\in\mathcal{H}$.
Proof
Item (i): the statement is valid since, due to Proposition 1(i), sequence $\left(x^{k}\right)_{k\in\mathds{N}}$ is bounded.
Item (ii): Take any $y\in M$. For all $k\in\mathds{N}$, we get
$\left\|x^{k}-y\right\|^{2}-\|y\|^{2}=\left\|x^{k}\right\|^{2}-2\left\langle y,x^{k}\right\rangle.$
Moreover, it follows from Proposition 1(ii) that $\lim\left(\left\|x^{k}-y\right\|^{2}-\|y\|^{2}\right)$ is well-defined.
Now, given two points $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, such that $x^{j_{k}}\rightharpoonup w^{\prime}$ and $x^{\ell_{k}}\rightharpoonup w^{\prime\prime}$, and considering Section 4 we have
$\lim\left\|x^{j_{k}}\right\|^{2}-2\left\langle y,w^{\prime}\right\rangle=\lim\left\|x^{\ell_{k}}\right\|^{2}-2\left\langle y,w^{\prime\prime}\right\rangle,$
and therefore
$\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\frac{1}{2}\left(\lim\left\|x^{j_{k}}\right\|^{2}-\lim\left\|x^{\ell_{k}}\right\|^{2}\right).$
Setting $\alpha$ as the right-hand side of the above equation yields that
$M\subset\left\{y\in\mathcal{H}\mid\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha\right\},$
as required.
Item (iii): In view of item (ii), if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$ then, for all $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$, there exists an $\alpha\in\mathds{R}$ such that for all $y\in\mathcal{H}$ it holds that $\left\langle y,w^{\prime}-w^{\prime\prime}\right\rangle=\alpha$. Consequently, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$ reduces to a singleton. Due to Proposition 2(i), $\left(x^{k}\right)_{k\in\mathds{N}}$ lies in a weakly compact set and therefore converges weakly.
Item (iv): Consider any $y\in\mathcal{H}$. Then, for all $k\in\mathds{N}$, we have
$\left\|x^{k}-y\right\|^{2}=\left\|x^{k}-\bar{x}\right\|^{2}+2\left\langle x^{k}-\bar{x},\bar{x}-y\right\rangle+\|\bar{x}-y\|^{2}.$
Using the hypothesis and Proposition 2 on the right-hand side of the above equation, we conclude the convergence of $\left(\left\|x^{k}-y\right\|\right)_{k\in\mathds{N}}$.
Theorem 4.2
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, $\left(x^{k}\right)_{k\in\mathds{N}}$ converges weakly to a point in $M$ if, and only if, $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$.
Proof
It holds trivially that Fejér* monotonicity together with weakly convergence implies that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$. On the other hand, suppose that $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and take $w^{\prime},w^{\prime\prime}\in\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}$. Thus, $w^{\prime},w^{\prime\prime}\in M$. Proposition 3(ii) affirms that $\left\langle w^{\prime},w^{\prime}-w^{\prime\prime}\right\rangle=\left\langle w^{\prime\prime},w^{\prime}-w^{\prime\prime}\right\rangle$ and hence, $w^{\prime}=w^{\prime\prime}$. The boundedness of $\left(x^{k}\right)_{k\in\mathds{N}}$, guaranteed by Proposition 1, completes the proof. ∎
We conclude this section by presenting a pair of results that characterizes the strong convergence of Fejér* monotone sequences.
Proposition 4
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then,
  1. (i)
    for all $w^{\prime},w^{\prime\prime}\in\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$, it holds that
    $M\subset\left\{y\in\mathcal{H}\mid\left\langle y-\frac{w^{\prime}+w^{\prime\prime}}{2},w^{\prime}-w^{\prime\prime}\right\rangle=0\right\};$
  2. (ii)
    if $\operatorname{cl}\left({\operatorname{aff}(M)}\right)=\mathcal{H}$, then $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}$ contains at most one point;
  3. (iii)
    the sequence converges strongly if there exist $x\in M,\left(\varepsilon_{k}\right)_{k\in\mathds{N}}\in\ell_{+}\cap\ell^{1}$, and $\rho\in(0,+\infty)$ such that for all $k\in\mathds{N}$ it holds that
    $\left\|x^{k+1}-x\right\|^{2}\leq\left\|x^{k}-x\right\|^{2}-\rho\left\|x^{k+1}-x^{k}\right\|+\varepsilon_{k}.$
Proof
It follows from Theorem 4.1 together with Proposition 3.7 in Combettes:2001a . ∎
Theorem 4.3
Let $\left(x^{k}\right)_{k\in\mathds{N}}$ be a Fejér* sequence with respect to a nonempty set $M\subset\mathcal{H}$. Then, the following statements are equivalent:
  1. (i)
    $\left(x^{k}\right)_{k\in\mathds{N}}$ converges strongly to a point in $M$;
  2. (ii)
    $\mathfrak{W}\left(x^{k}\right)_{k\in\mathds{N}}\subset M$ and $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}\neq\varnothing$;
  3. (iii)
    $\mathfrak{S}\left(x^{k}\right)_{k\in\mathds{N}}\cap M\neq\varnothing$.
Proof
This is a direct result of combining Theorem 4.1 and Theorem 3.11 in Combettes:2001a . ∎
We emphasize that new characterizations of Fejér* monotonicity could be revealed with further analysis, suggesting a direction for future research.
5 Concluding remarks
In this work we further expand the understanding of Fejér* monotonicity. We provide a characterization of Fejér* monotonicity in terms of the convergence of scalar sequences and inner product sequences to the closure of the convex hull of the underlying set. We provide an insightful example that shows that Fejér* monotonicity not always can be extended to the closure of the convex hull of the underlying set. It was also established connections between Fejér* monotonicity and quasi-Fejér monotonicity. The results presented in this paper could shed a light on the analysis of the (weak) convergence of algorithms that comply with Fejér* monotonicity.
Acknowledgements.
RB was partially supported by Brazilian agency Fundação de Amparo à Pesquisa do Estado do Rio de Janeiro (Grant E-26/201.345/2021); YBC was partially supported by the USA agency National Science Foundation (Grant DMS-2307328); AAR was partially supported by Brazilian agency Conselho Nacional de Desenvolvimento Científico e Tecnológico – CNPq (Grant 307987/2023-0); LRS was also partially funded by CNPq (Grant 310571/2023-5).